# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

### v2

@@ VIII-D Proof of Theorem 2
and a point $x^{*}\in\mathbb{R}^{n}$. We say $(u_{s}(x^{*}),\lambda(x^{*}))$ is a regular minimizer if the LICQ is satisfied at $u_{s}(x^{*})$, and $(u_{s}(x^{*}),\lambda(x^{*}))$ satisfies the KKT condition and the strong second order sufficient condition (SSOSC). The point $(u,\lambda)$ satisfies the SSOSC if Hessian matrix of the Lagrangian of 66 is positive definite at this point. According to [25, Theorem 2.3.3], there exists a locally Lipschitz continuous map $u_{s}(x)$ around a neighborhood of $x^{*}$ such that $u_{s}(x)\in\mathbb{R}^{m}$ is a minimizer for the non-linear optimization problem 66, if $f$ and $g$ are twice continuously differentiable in $(u,x)$, and $u_{s}(x^{*})$ is a regular minimizer with multiplier $\lambda(x^{*})$. In what follows, we establish these conditions for the QCQP in 43.
LICQ property under Assumption 3 asserts that for any solution of the QCQP $u_{s}(x^{*})$, there exists a unique $\lambda^{*}$ that satisfies the KKT condition at $(u_{s}(x^{*}),\lambda^{*})$. Next, we show that the SSOSC holds for any choice of $\lambda^{*}$, specifically $\lambda^{*}$ that satisfies the KKT condition. Consider the Lagrangian of the QCQP for either linear or quadratic input constraints:
$\begin{split}L(u,\lambda)=&u^{T}u+2u_{n}(x)u+u_{n}(x)^{T}u_{n}(x)\\ &+\lambda_{1}^{T}(C(x)u+b(x))\\ &+\lambda_{2}^{T}(u^{T}Q_{u}u-u_{max}^{2})+\lambda_{3}^{T}(Au-b)\end{split}$
Then the inequality $v^{T}\nabla_{uu}^{2}L(u,\lambda)v=v^{T}v+\lambda_{2}v^{T}Q_{u}v>0$ for all $v\in\mathbb{R}^{m}$. This proves that any solution of the QCQP $u_{s}(x^{*})$ satisfies the SSOSC with the corresponding $\lambda^{*}$. Furthermore, a solution of the QCQP $u_{s}(x)$ exists for all $x$ due to the lower bound on $r(x)$ in 48, consequently establishing the local Lipschitz continuity of $u_{s}(x)$.

### v3

@@ VIII-E Proof of Theorem 2
and a point $x^{*}\in\mathbb{R}^{n}$. We say $(u_{s}(x^{*}),\lambda(x^{*}))$ is a regular minimizer if the LICQ is satisfied at $u_{s}(x^{*})$, and $(u_{s}(x^{*}),\lambda(x^{*}))$ satisfies the KKT condition and the strong second order sufficient condition (SSOSC). The point $(u,\lambda)$ satisfies the SSOSC if Hessian matrix of the Lagrangian of 71 is positive definite at this point. According to [9, Theorem 2.3.3], there exists a locally Lipschitz continuous map $u_{s}(x)$ around a neighborhood of $x^{*}$ such that $u_{s}(x)\in\mathbb{R}^{m}$ is a minimizer for the non-linear optimization problem 71, if $f$ and $g$ are twice continuously differentiable in $(u,x)$, and $u_{s}(x^{*})$ is a regular minimizer with multiplier $\lambda(x^{*})$. In what follows, we establish these conditions for the QP in 44.
LICQ property under Assumption 3 asserts that for any solution of the QP $u_{s}(x^{*})$, there exists a unique $\lambda^{*}$ that satisfies the KKT condition at $(u_{s}(x^{*}),\lambda^{*})$. Next, we show that the SSOSC holds for any choice of $\lambda^{*}$, specifically $\lambda^{*}$ that satisfies the KKT condition. Consider the Lagrangian of the QP for either linear or quadratic input constraints:
$\begin{split}L(u,\lambda)=&u^{\top}u+2u_{n}(x)u+u_{n}(x)^{\top}u_{n}(x)\\ &+\lambda_{1}^{\top}(C(x)u+b(x))\\ &+\lambda_{2}^{\top}(u^{\top}Q_{u}u-u_{max}^{2})+\lambda_{3}^{\top}(Au-b)\end{split}$
Then the inequality $v^{\top}\nabla_{uu}^{2}L(u,\lambda)v=v^{\top}v+\lambda_{2}v^{\top}Q_{u}v>0$ for all $v\in\mathbb{R}^{m}$. This proves that any solution of the QP $u_{s}(x^{*})$ satisfies the SSOSC with the corresponding $\lambda^{*}$. Furthermore, a solution of the QP $u_{s}(x)$ exists for all $x$ due to the lower bound on $r(x)$ in 49, consequently establishing the local Lipschitz continuity of $u_{s}(x)$.